Use the table of areas under the standard normal curve to find the probability that a z-score from the standard normal distribution will lie within the interval. (Round your answer to four decimal places.)

0 ≤ z ≤ 1.5

−0.8 ≤ z ≤ 0

−0.8 ≤ z ≤ −0.6

−1.6 ≤ z ≤ 2.4
Answer #1

solution

(A)P( 0≤ Z ≤1.5 )

= P(Z ≤ 1.5) - P(Z ≤0 )

Using z table,  

= 0.9332-0.5  

area=0.4332

(B)

P( −0.8 ≤ z ≤ 0)

= P(Z ≤ 0) - P(Z ≤-0.8 )

Using z table,  

= 0.5-0.2119

probability=0.2881

(C)P(−0.8 ≤ z ≤ −0.6 )

= P(Z ≤-0.6 ) - P(Z ≤ -0.8)

Using z table,  

= 0.2743-  0.2119

probability=0.0624

(D)

P(−1.6 ≤ z ≤ 2.4)

= P(Z ≤2.4 ) - P(Z ≤ -1.6)

Using z table,  

= 0.9918-0.0548

probability=0.9370